# GECAM Observations of the Galactic Magnetar SGR J1935+2154 during the 2021 and 2022 Burst Active Episodes. I. Burst Catalog

### v3

@@ 3 Catalog Analysis
As shown above, we show light curves of 9 bursts in Fig. 4 to display the variety of these bursts. The trigger time of each burst is listed in Tables 3 and 4. SGR J1935+2154 bursts consist of a single pulse (such as ID: 122, 138, and 161 of GECAM-B; ID: 3 and 42 of GECAM-C), multiple pulses (ID: 27 of GECAM-B; ID: 59 of GECAM-C), and intermediate flares (ID: 126 of GECAM-B; ID: 23 of GECAM-C). The X-ray burst associated with the fast radio burst FRB 20221014 is ID: 138 of GECAM-B, and ID: 42 of GECAM-C.
@@ 3.1 Burst Activity
We calculate the effective observation time  [^3] GECAM-B orbits at low latitudes ($<30^{\circ}N/S$), while GECAM-C could cover nearly $90^{\circ}N/S$. We only calculate the observation time of GECAM-C at low latitudes in this paper since the background undergoes too violent changes to detect magnetar bursts when GECAM-C is across high latitudes. of GECAM-B/C individually and estimate the daily burst rate of GECAM-B/C, respectively (see Fig. 5). Because the effective detectable energy range of GECAM-B changed in the HG channel (¿40 keV after October 2022), weaker bursts may not be detected compared to GECAM-C observations. Hence, given that the burst activity of GECAM-B is lower than that of GECAM-C (¿15 keV), the daily burst rate estimated using GECAM-C data may be more reliable for assessing burst activity compared to using GECAM-B data after October 2022. A detailed study on the evolution of the effective detectable energy range change of GECAM-B will be reported in Paper II.
@@ 3.2 Period Search
To explore the potential periodicity behavior in the burst history of SGR J1935+2154, we use the daily burst rate data from GECAM-B (before October 2022) and GECAM-C (after October 2022) with the Lomb-Scargle Periodogram for searching periodicity (Lomb, 1976; Scargle, 1982; VanderPlas, 2018). The periodogram results are shown in Fig 6. The most significant peak of the Lomb-Scargle periodogram is 134.63 day (False Alarm Probability: 0.001). The peak around the 50 day are likely caused by observation windows or bursts gap (Xie et al., 2022).
To uncover the active period of SGR J1935+2154, the start time of all bursts are folded into different phases using period folding method (CHIME/FRB Collaboration et al., 2020b; Zhang et al., 2021; Xie et al., 2022),
@@
where $\phi$ is the folded phase, $T$ is the burst start time, $T_{0}$ is the phase start point (MJD: 59215; UTC: 2021-01-01), $P$ is a given period, and the floor is a function that returns the floor of the input number. As shown in Fig 7, the duty circle of SGR J1935+2154 bursts is 80%. The error could be estimated to be 20 days with
@@
where $P$ is the period and $W_{\mathrm{active}}$ is the number of active days. Accordingly, the most probable period of SGR J1935+2154 from January 2021 to December 2022 is $134.63\pm 20$ days, consistent with previous work (Xie et al., 2022, evaluated from July 2014 to January 2022) and the burst history over these two years could be grouped into 4 active episodes (see Fig. 5).
@@ 3.3 Burst Duration
As previous studies (Lin et al., 2013), the magnetar burst duration $T_{\mathrm{bb}}$ is calculated with the Bayesian blocks method (Scargle et al., 2013). We measure the duration for each burst using the sliced event data of 10 s, including both pre- and post-burst trigger time intervals (from T0-10 s to T0+10 s, where T0 is the burst trigger time). The blocks with a duration longer than 6 s are treated as background since the spin period of SGR J1935+2154 is just approximately 3.24 s. The background count rate is estimated by fitting a first-order polynomial to the background blocks. A block with a significance ($S/\sqrt{B}$, where $S$ is the net counts and $B$ is the background counts) larger than $5\sigma$ is considered as part of the burst region. As an example, the blocks and block edges of a burst are illustrated in Fig. 4. However, some bursts with multiple pulses (e.g. burst #59 of GECAM-C in Fig. 4) have at least two subsequent blocks, along with quiescent intervals. In such cases, the entire burst duration is calculated from the start of the first burst block to the end of the last block. Due to the different effective detectable energy ranges between GECAM-B and GECAM-C, 14 bursts observed by both GECAM-B&C exhibit different durations (see Fig. 8 or see an example of X-ray burst #138 of GECAM-B and #42 of GECAM-C in Fig. 4 associated with FRB 221014).
The duration ($T_{\mathrm{bb}}$) of each burst observed by GECAM-B/C is listed in Tables 3 and 4. The burst duration distributions of GECAM-B/C are shown in Fig. 9. These distributions are well fitted with lognormal functions, as illustrated in Table 5. We conduct the one-sample Kolmogorov-Smirnov test to compare the distribution of GECAM-B/C against a lognormal distribution and show the goodness of fit in Table 5. The results show that the distribution of burst duration detected by GECAM from SGR J1935+2154 exhibits a lognormal distribution, which is consistent with previous work (Lin et al., 2020b, a; Cai et al., 2022a; Rehan & Ibrahim, 2023, 2024).
@@
GECAM-B 1st episode $45.50_{-32.56}^{+114.46}$ 0.18 0.64
2nd episode $65.77_{-43.12}^{+125.22}$ 0.06 0.98
3rd episode $120.78_{-79.38}^{+231.59}$ 0.11 0.57
4th episode $86.30_{-60.65}^{+204.10}$ 0.11 0.58
GECAM-C 4th episode $83.37_{-66.35}^{+324.95}$ 0.07 0.69
@@ 3.4 Burst Waiting Time
Fig. 10 shows the distribution of waiting time measured by GECAM-B/C. These distributions could be well fitted with lognormal functions, resulting in a mean value of $338.84_{-212.95}^{+573.16}$ seconds for GECAM-B and a mean value of $79.43_{-51.89}^{+149.65}$ seconds for GECAM-C. When comparing to a lognormal distribution, the Kolmogorov-Smirnov tests yield a statistic of 0.104 and a p-value of 0.672 for GECAM-B, and a statistic of 0.112 and a p-value of 0.929 for GECAM-C. This lognormal distribution behavior is similar to that observed in previous studies (Cai et al., 2022a; Xie et al., 2024). Since the relatively short continuous observation time intervals of GECAM-C for observing SGR J1935+2154 (only dozens of minutes) compared to GECAM-B, it is difficult to measure the longer waiting times between more successive bursts. Therefore, the distribution mean value of GECAM-C is smaller than that of GECAM-B.
Figure 10: The left panel is the waiting time (time interval between two successive bursts observed by GECAM-B/C) of the X-ray bursts v.s their occurring times, where the time of FRB 20221014 is represented by the green dashed line. The right panel is the distribution of waiting time. The horizontal dotted lines represent the best-fit lognormal function with a mean of $338.84_{-212.95}^{+573.16}$ seconds for GECAM-B and $79.43_{-51.89}^{+149.65}$ seconds for GECAM-C, respectively.
@@ 3.5 Burst Hardness Ratio
where $C_{\mathrm{i}}$ and $B_{\mathrm{i}}$ represent the total counts and background counts in the burst duration, respectively. The hardness ratio is the ratio of the net counts in different energy ranges.
Since changes in the effective detectable energy range of GECAM-B over the two years, we measure the hardness ratio of all bursts within the same effective range of 50-200/40-50 keV for GECAM-B. The hardness ratio of GECAM-C is calculated in the range of 30-200/15-30 keV. The hardness ratio of the four active episodes for GECAM-B data yields mean values of 2.70, 2.30, 1.87, and 1.48, respectively. It presents that the evolution of the burst hardness tends to be softer across the observation time, as shown in Fig. 11. The burst hardness is more concentrated after October 2022 when FRB 20221014 happened. More detailed spectral properties and evolution will be reported in Paper II.
@@ 4 Summary and Discussions
In this paper, we perform a targeted sub-threshold search GECAM daily EVT data for SGR J1935+2154 bursts during the year of 2021 and 2022. We design 5 spectral templates to target search and find 164 X-ray bursts from GECAM-B and 97 bursts from GECAM-C (refer to Tables 3 and 4). Among the templates, the CPL, OTTB, and BB cover a higher proportion of them compared to others (refer to Table 2), providing more accuracy in location (see Fig. 2). Therefore, a simple power-law function and exponential cutoff function with softer parameter settings might be preferred for searching magnetar X-ray bursts.
@@
The burst activity is evaluated by estimating the daily burst rate data, as illustrated in Fig. 5. We employ the Lomb-Scargle method to investigate potential periodic behavior, resulting in the most likely period of $134.63\pm 20$ days (refer to Fig. 6). Therefore, the burst history over these two years could be divided into 4 active episodes (see Fig. 5). The period of $134.63\pm 20$ days with 80% duty circle is well consistent with previous study (Xie et al., 2022, evaluated from July 2014 to January 2022). Unlike other study periods ($\sim 231$ days evaluated from 2014 to 2020 or $\sim 238$ days evaluated from July 2014 to October 2021 in Grossan, 2021; Zou et al., 2021, respectively), a shorter period is found in this study, possibly indicating an increased activity of X-ray bursts from SGR J1935+2154 after the year of 2021.
@@
We perform extensive studies on the statistical characteristics of these X-ray bursts, including the burst duration, waiting time, and hardness ratio. The duration follows a lognormal distribution for GECAM-B/C, respectively (see Fig. 9 and Table 5). Due to the varying effective detectable energy ranges of GECAM-B and GECAM-C, some bursts detected by both satellites may show different durations (see Fig. 4 for examples). In such cases, the duration measured by GECAM-C is more informative across the entire energy ranges when both satellites detect the same burst. The waiting time also conforms to a lognormal function with mean values of $338.84_{-212.95}^{+573.16}$ seconds and $79.43_{-51.89}^{+149.65}$ seconds for GECAM-B/C, respectively (refer to Fig. 10). Because the relatively short continuous observation time intervals for GECAM-C observing SGR J1935+2154 compared to GECAM-B, the distribution mean value of GECAM-C is smaller than that of GECAM-B. Such lognormal distribution behavior is similar to that observed in previous studies (Cai et al., 2022a; Xie et al., 2024). The hardness ratio of X-ray bursts tends to become softer and more concentrated over the course of the observation time spanning these two years (see Fig. 11).
@@
These statistical results indicate that the X-ray bursts from SGR J1935+2154 show increased activity and exhibit a softer spectrum over the two years. Similar to FRB 200428, which occurred during the most active period (April-May 2020) preceding 2020, the FRB 20221014 are detected in the final and most active episode (October-December 2022). In view of their millisecond duration and high energy releases, FRBs are widely suggested to originate from the violent activities of compact objects, in particular, magnetars (Popov & Postnov, 2010; Kulkarni et al., 2014; Katz, 2016; Connor et al., 2016; Cordes & Wasserman, 2016; Lyutikov, 2017). And the amounts of energy released by the bursts and outbursts of Galactic magnetars, from SGRs or AXPs, indicate that X-ray bursts probably correspond to sudden and local (not global) catastrophic events. In that case, FRB emission may result from such burst interactions (e.g., the collision between different Alfven waves or different explosion outflows) in a global activity of a magnetar and then a detected FRB be associated with an X-ray burst (Yang & Zhang, 2021; Xie et al., 2024).

### v5

@@ 3 Catalog Analysis
As shown above, we show light curves of 9 bursts in Fig. 4 to display the variety of these bursts. The trigger time of each burst is listed in Tables 3 and 4. SGR J1935+2154 bursts consist of a single peak (such as ID: 122, 138, and 161 of GECAM-B; ID: 3 and 42 of GECAM-C), multiple peaks (ID: 27 of GECAM-B; ID: 59 of GECAM-C), and intermediate bursts (ID: 126 of GECAM-B; ID: 23 of GECAM-C). The X-ray burst associated with the radio burst FRB 20221014 is ID: 138 of GECAM-B and ID: 42 of GECAM-C.
@@ 3.1 Burst Activity
We calculate the effective observation time [^8] GECAM-B orbits at low latitudes ($<30^{\circ}N/S$), while GECAM-C could cover nearly $90^{\circ}N/S$. We only calculate the observation time of GECAM-C at low latitudes in this paper since the background undergoes too violent changes to detect magnetar bursts when GECAM-C is across high latitudes. of GECAM-B/C individually and estimate the daily burst rate of GECAM-B/C, respectively (see Fig. 5). Because the effective detectable energy range of GECAM-B changed in the HG channel ($>$40 keV after October 2022), weaker bursts may not be detected compared to GECAM-C observations. Hence, given that the burst activity of GECAM-B is lower than that of GECAM-C ($>$15 keV), the daily burst rate estimated using GECAM-C data may be more reliable for assessing burst activity compared to using GECAM-B data after October 2022. A detailed study on the evolution of the effective detectable energy range change of GECAM-B will be reported in Paper II.
@@ 3.2 Period Search
To explore the potential periodicity behavior in the burst history of SGR J1935+2154, we use the daily burst rate data of GECAM-B (before October 2022) and GECAM-C (after October 2022) to search periodicity by the Lomb-Scargle Periodogram method (Lomb, 1976; Scargle, 1982; VanderPlas, 2018). The periodogram results are shown in Fig 6. The most significant peak of the Lomb-Scargle periodogram is 134.63 day (false alarm probability: 0.001, significance level $\sim 3\sigma$). The peaks around the 50 day are likely caused by observation windows or bursts gap (Xie et al., 2022).
To uncover the burst activity of SGR J1935+2154 within a cycle, the start time of all bursts is folded into different phases at a given period $P=134.63$ day through the period folding method (CHIME/FRB Collaboration et al., 2020b; Zhang et al., 2021; Xie et al., 2022),
@@
where $\phi$ is the folded phase, $T$ is the burst start time, $T_{0}$ is the phase start point (MJD: 59215; UTC: 2021-01-01), and the floor is a function that returns the floor of the input number. As shown in Fig 7, the active day within a cycle (duty cycle) of SGR J1935+2154 bursts is 80%. The error of the period $P=134.63$ could be estimated to be 20 days with (Chime/Frb Collaboration et al., 2020),
@@
where $W_{\mathrm{active}}$ is the active days in one cycle and $T_{\mathrm{span}}$ is the longest time separation between burst arrival times. Accordingly, the most probable period of SGR J1935+2154 from January 2021 to December 2022 is $134.63\pm 20$ days, consistent with previous work (Xie et al., 2022, evaluated from July 2014 to January 2022) and the burst history over these two years could be grouped into 4 active episodes (see Fig. 5).
@@ 3.3 Burst Duration
As previous studies (Lin et al., 2013), the magnetar burst duration $T_{\mathrm{bb}}$ is calculated with the Bayesian blocks algorithm (Scargle et al., 2013). We measure the duration for each burst using the sliced event data of 10 s, including both pre- and post-burst trigger time intervals (from T0-10 s to T0+10 s, where T0 is the burst trigger time). The blocks with a duration longer than 6 s (at least twice the spin period of SGR J1935+2154) are treated as background. The background count rate is estimated by fitting a first-order polynomial to the background blocks. A block with a significance larger than $5\sigma$ is considered to be part of the burst region. The first block with a significance higher than $5\sigma$ is marked as the burst start time (see Tables 3 and 4). As an example, the blocks and block edges of a burst are illustrated in Fig. 4. However, some bursts with multiple peaks (e.g., burst #59 of GECAM-C in Fig. 4) have at least two subsequent blocks, along with quiescent intervals. In such cases, the duration of the entire burst is calculated from the start of the first burst block to the end of the last block. Due to the different effective detectable energy ranges between GECAM-B and GECAM-C, 14 bursts observed by both GECAM-B&C exhibit different durations (see Fig. 8 or see an example of X-ray burst #138 of GECAM-B and #42 of GECAM-C in Fig. 4 associated with FRB 221014).
The duration ($T_{\mathrm{bb}}$) of each burst observed by GECAM-B/C is listed in Tables 3 and 4. The burst duration distributions of GECAM-B/C are shown in Fig. 9. These distributions are well fitted with lognormal functions, as illustrated in Table 5. We perform the one-sample Kolmogorov-Smirnov test to compare the distribution of GECAM-B/C against a lognormal distribution and show the goodness of fit in Table 5. The mean values of burst duration for the four active episodes are $47.34_{-34.35}^{+125.19}$ ms for the first episode, $70.21_{-45.05}^{+125.69}$ ms for the second episode, $120.75_{-79.32}^{+231.22}$ ms for the third episode, and $85.51_{-60.77}^{+210.03}$/$83.33_{-66.32}^{+324.99}$ ms (given by GECAM-B/C data) for the fourth active episodes. The full sample gives the mean value $83.48_{-61.57}^{+234.64}$ ms. The results show that the distribution of burst duration detected by GECAM from SGR J1935+2154 exhibits a lognormal distribution, which is consistent with previous work (Lin et al., 2020b, a; Cai et al., 2022a; Rehan & Ibrahim, 2023, 2024).
@@
GECAM-B 1st episode $47.34_{-34.35}^{+125.19}$ 0.2 0.60
2nd episode $70.21_{-45.05}^{+125.69}$ 0.07 0.95
3rd episode $120.75_{-79.32}^{+231.22}$ 0.11 0.57
4th episode $85.51_{-60.77}^{+210.03}$ 0.13 0.48
GECAM-C 4th episode $83.33_{-66.32}^{+324.99}$ 0.07 0.69
@@ 3.4 Burst Waiting Time
Fig. 10 shows the distribution of waiting time measured by GECAM-B/C. These distributions could be well fitted with lognormal functions, resulting in a mean value of $338.84_{-212.95}^{+573.16}$ seconds for GECAM-B and a mean value of $79.43_{-51.89}^{+149.65}$ seconds for GECAM-C. Compared to a lognormal distribution, the Kolmogorov-Smirnov tests yield a statistic of 0.104 and a p-value of 0.672 for GECAM-B, and a statistic of 0.112 and a p-value of 0.929 for GECAM-C. This lognormal distribution behavior is similar to that observed in previous studies (Cai et al., 2022a; Xie et al., 2024). Since the relatively short continuous observation time intervals of GECAM-C for observing SGR J1935+2154 (only dozens of minutes) compared to GECAM-B, it is difficult to measure the longer waiting times between more successive bursts. Therefore, the distribution’s mean value of GECAM-C is smaller than that of GECAM-B.
Figure 10: The left panel is the waiting time (time interval between two successive bursts observed by GECAM-B/C) of the X-ray bursts v.s their occurring times, where the time of the FRB 20221014 is represented by the green dashed line. The right panel is the distribution of waiting time. The horizontal dotted lines represent the best-fit lognormal function with a mean of $338.84_{-212.95}^{+573.16}$ seconds for GECAM-B and $79.43_{-51.89}^{+149.65}$ seconds for GECAM-C, respectively.
@@ 3.5 Burst Hardness Ratio
where $C_{\mathrm{i}}$ and $B_{\mathrm{i}}$ represent the total counts and background counts in the burst duration, respectively. The hardness ratio is the ratio of net counts in different energy ranges.
Since changes in the effective detectable energy range of GECAM-B over the two years, we measure the hardness ratio of all bursts within the same effective range of 50-200/40-50 keV for GECAM-B. The hardness ratio of GECAM-C is calculated in the range of 30-200/15-30 keV. The hardness ratio of the four active episodes for the GECAM-B data yields mean values of $2.72\pm 1.70$, $2.25\pm 1.20$, $1.87\pm 0.63$, and $1.48\pm 0.47$, respectively. The hardness ratio of the last episode for the GECAM-C data yields a mean value of $0.76\pm 0.33$. It presents that the evolution of the burst hardness tends to be softer through the observation time, as shown in Fig. 11. More detailed spectral properties and evolution will be reported in Paper II.
@@ 4 Summary and Discussions
In this paper, we perform a targeted sub-threshold search GECAM daily EVT data for SGR J1935+2154 bursts during the year of 2021 and 2022. We design five spectral templates to target search and find 159 X-ray bursts from GECAM-B and 97 bursts from GECAM-C (refer to Tables 3 and 4). Among the templates, CPL, OTTB, and BB cover a higher proportion of them compared to others (refer to Table 2), providing more accuracy in location (see Fig. 2). Therefore, a simple power-law function and exponential cutoff function with softer parameter settings might be preferred for searching magnetar X-ray bursts.
@@
The burst activity is evaluated by estimating the daily burst rate data, as illustrated in Fig. 5. We employ the Lomb-Scargle method to investigate potential periodic behavior, which results in the most likely period of $134.63\pm 20$ days (refer to Fig. 6). Therefore, the burst history over these two years could be divided into four active episodes (see Fig. 5). The period of $134.63\pm 20$ days with 80% duty cycle is well consistent with previous study (Xie et al., 2022, evaluated from July 2014 to January 2022). Unlike other periods derived in previous work ($\sim 231$ days evaluated from 2014 to 2020 or $\sim 238$ days evaluated from July 2014 to October 2021 in Grossan, 2021; Zou et al., 2021, respectively), a shorter period is found in this study, possibly indicating an increased activity of X-ray bursts from SGR J1935+2154 after the year of 2021.
Table 6: Comparison of the burst duration of SGR 1935+2154 burst episodes since its discovery in 2014, as Observed by the Fermi/GBM and GECAM-B/C.
Episode Instrument Number of bursts DurationaaThe lognormal mean value of burst duration.
Jul 2014 GBM 3
Feb-Mar 2015 GBM 24 $78.00_{-14.00}^{+17.00}$
May-Jun 2016 GBM 42 $72.00_{-6.00}^{+7.00}$
Jun-Jul 2016 GBM 54 $128.00_{-10.00}^{+11.00}$
Nov 2019 GBM 22 $121.00_{-33.00}^{+45.00}$
Apr-May 2020 GBM 151 $182.00_{-19.00}^{+22.00}$
Jan 2021 GECAM-B 14 $47.34_{-34.35}^{+125.19}$
Sep 2021 GBM 79 $77.34\pm 1.03$
GECAM-B 57 $70.21_{-45.05}^{+125.69}$
Jan 2022 GBM 112 $97.03\pm 1.03$
GECAM-B 46 $120.75_{-79.32}^{+231.22}$
Oct-Dec 2022 GECAM-B 42 $85.51_{-60.77}^{+210.03}$
GECAM-C 97 $83.33_{-66.32}^{+324.99}$
@@
We perform extensive studies on the statistical characteristics of these X-ray bursts, including the burst duration, waiting time, and hardness ratio. The duration follows a lognormal distribution for GECAM-B/C, respectively (see Fig. 9 and Table 5). Due to the varying effective detectable energy ranges of GECAM-B and GECAM-C, some bursts detected by both satellites may show different durations (see Fig. 4 for examples). In such cases, the duration measured by GECAM-C is more informative across the entire energy range when both satellites detect the same burst. The burst duration of SGR J1935+2154 is around 100 ms, see Table 6 for comparison among different burst episodes. The waiting time also conforms to a lognormal function with mean values of $338.84_{-212.95}^{+573.16}$ seconds and $79.43_{-51.89}^{+149.65}$ seconds for GECAM-B/C, respectively (refer to Fig. 10). Because of the relatively short continuous observation time intervals for GECAM-C observing SGR J1935+2154 compared to GECAM-B, the distribution’s mean value of GECAM-C is smaller than that of GECAM-B. Such lognormal distribution behavior is similar to that observed in previous studies (Cai et al., 2022a; Xie et al., 2024). The hardness ratio of X-ray bursts tends to become softer over the course of the observation time spanning these two years (see Fig. 11).
@@
These statistical results indicate that the X-ray bursts from SGR J1935+2154 show increased activity, as observed by Fermi/GBM (e.g., Rehan & Ibrahim, 2024). The magnetar bursts also exhibit a softer spectrum over the two-year period, as presented by the temporal evolution of the hardness ratio and a detailed spectral analysis will follow. Similar to FRB 200428, which occurred during the most active episode (April-May 2020) preceding 2020, the radio burst FRB 20221014 are detected in the final and most active episode (October-December 2022). In view of their millisecond duration and high energy releases, FRBs are widely suggested to originate from the violent activities of compact objects, in particular magnetars (Popov & Postnov, 2010; Kulkarni et al., 2014; Katz, 2016; Connor et al., 2016; Cordes & Wasserman, 2016; Lyutikov, 2017). Leading to such burst activities may correspond to some sudden changes in the magnetic configuration of the magnetars (e.g., starquakes, magnetic field reconnections Thompson & Duncan, 1995; Thompson & Blaes, 1998; Jones, 2003; Levin & Lyutikov, 2012). In that case, FRB emission may result from such burst interactions (e.g., the collision between different Alfven waves or different explosion outflows) in a global activity of a magnetar and then a detected FRB be associated with an X-ray burst (Yang & Zhang, 2021; Xie et al., 2024).